OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 5 and √2

OpenStudy (campbell_st):

 use pythagoras' theorem \[x^2 = 5^2 + (\sqrt{2})^2\] just evaluate for the answer

OpenStudy (anonymous):

square both numbers add the squares take the square root of the sum of the squares

OpenStudy (anonymous):

2 √3?

OpenStudy (campbell_st):

umm.... not quite... certainly root 3 but the number at the front is incorrect

OpenStudy (anonymous):

3 √3?

OpenStudy (campbell_st):

thats it...